A gas of volume 2 cubic meters at 27 degree Celsius  is heated to 327 degree Celsius at constant pressure. What is its new volume? 

Grade:11

1 Answers

Khimraj
7 years ago
At constant pressure for gas the volume is directly proportional to absolute temperature soV1/V2 = T1/T22/V2 = 300K/600KSo V2 = 4 cubic meters